Calculate the discriminant to determine the number of real roots.
y = x2 + 3x + 9
How many real roots does the equation have?
one real root
two real roots
no real roots
no solution to the equation

Answer by stanbon(75887)

Calculate the discriminant to determine the number of real roots.
y = x2 + 3x + 9
D = b^2-4ac
D = 3^2 - 4*1*9 = 9-36 = -27
-----
Since D is negative the equation has no Real roots
-------------------------------------------------------
Cheers,
Stan H.
-------------
How many real roots does the equation have?
one real root
two real roots
no real roots
no solution to the equation